### v2

# GAI_arXiv.bib

Abstract
This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, it proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of innovation. The core of the GAI framework lies in an architecture that dynamically processes the internal states of agents and a dialogue scheme specifically tailored to facilitate analogy-driven innovation. The framework’s functionality is evaluated using Dyson’s invention of the bladeless fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through a set of fictional technical documents. The experimental results demonstrate that models with internal states significantly outperformed those without, achieving higher average scores and lower variance. Notably, the model with five heterogeneous agents equipped with internal states successfully replicated the key ideas underlying the Dyson’s invention. This indicates that the internal state enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
Keywords: generative AI, Large language models, Multi-Agent Systems, Innovation, Patent
I Introduction
Generative agents, including LLM agents, can engage in reflections for decision-making or analysis by breaking down complex tasks and conducting systematic reasoning. LLM-based Multi-agent System (LLM-MAS) extends this capability by adding collective reasoning functions, with applications already emerging in fields such as software development and social simulations ([chan2023chatevalbetterllmbasedevaluators, guo2024largelanguagemodelbased, hong2024metagptmetaprogrammingmultiagent, park2023generativeagentsinteractivesimulacra, qian2024chatdevcommunicativeagentssoftware, sun2024llmbasedmultiagentreinforcementlearning, Xiong_2023, zhou2024sotopiainteractiveevaluationsocial]). One particularly noteworthy area is the application of LLM-MAS to scientific discussions. In some studies, generative agents have been used to address tasks including mathematical problems and medical debates, with interesting findings such as agents correcting their mistakes through dialogue ([tang2024medagentslargelanguagemodels, du2023improvingfactualityreasoninglanguage]). Applying these capabilities of generative agents to discussions based on technical documents, such as patents, could potentially replicate the collective interactions that human teams engage in during the innovation process.
Efforts to analyze the innovation process have historically included observational experiments in actual corporate environments, but these have faced significant limitations due to low reproducibility and difficulties in manipulating experimental conditions. As a result, the effects of variables such as organizational structure, communication methods, and degree of diversity on the innovation process remain not fully understood. If LLM-MAS can replicate the innovation process, it would enable repeated simulations with adjustable conditions under high reproducibility.
Furthermore, it is conceivable that artificial intelligence technologies, including LLMs, will increasingly be applied in real-world corporate innovation settings, creating a new domain of practice that could be described as “AI-driven innovation.” If LLM-MAS can simulate the innovation process, it would also be possible to pioneer this new domain by integrating human teams with generative agent teams, demonstrating new possibilities for AI-human collaboration in driving innovation.
The objective of this study is to examine whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, the study specifically focuses on analogy-driven innovation and proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate this process.
The core of the GAI framework lies in an architecture that dynamically processes the internal states of agents and a dialogue scheme specifically designed to facilitate analogy-driven innovation. In the former, agents generate and refine ideas based on intrinsic motivation and critical introspection, leveraging previously reviewed documents and interactions with others to form their current thoughts, which serve as the basis for their contributions to the dialogue. In the latter, the scheme incorporates insights from management science and cognitive science to replicate the process of identifying functional similarities and mechanical differences, enabling the transfer of solutions across diverse technical domains.
To evaluate the functionality of the framework, this study uses Dyson’s invention of the bladeless household fan as a case study. It examines the extent to which analogy-driven innovations can be replicated by the proposed framework using a set of fictional technical documents.
II Proposed framework
II-A Agent architecture
The architecture of each agent consists of two primary modules driven by an LLM: a memory module and an internal state module (Figure 1). The memory module records the contents of reviewed technical documents, past conversations with others, and the agent’s prior reflections, all timestamped for reference. These records are then organized into a structured memory, which is input into the internal state module. The internal state module processes newly acquired technical information and statements from others during the dialogue, drawing on both recent thoughts and memories retrieved from the memory module. Based on this, the internal state module forms and updates the agent’s current thoughts, which serve as the foundation for its future reflections and statements.
Figure 1: Agent architecture and organizational structure
When forming and updating current thoughts, the internal state module undergoes two processes: idea generation and introspection.
During the idea generation process, the module generates multiple ideas for reflections on a technological domain based on the memories received from the memory module and the agent’s thoughts at that time. Subsequently, the internal state module evaluates these ideas on a scale of 1 to 10 for each of three criteria—Novelty, Importance, and Consensus—using a language model and selects ideas that maximize an intrinsic reward calculated as a weighted sum of these criteria:
$\begin{split}\max_{x_{ij}}r_{ij}&\equiv\alpha_{i}\,\text{Novelty}(x_{ij})+\beta_{i}\,\text{Importance}(x_{ij})\\ &+\gamma_{i}\,\text{Consensus}(x_{ij}),\end{split}$ (1)
where $r_{ij}$ represents the intrinsic reward of agent $i$, $x_{ij}$ is its idea on object $j$, and $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ respectively represent the weights for the 1-10 evaluation on each criterion, which are denoted by Novelty$(x_{ij})$, Importance$(x_{ij})$, and Consensus$(x_{ij})$ in Equation (1). The Novelty criterion assesses the degree to which an idea introduces new content that has not been present in prior conversations and reflections. The Importance criterion evaluates how significant the idea is in light of previous discussions and reflections. The Consensus criterion measures the extent to which the idea aligns with the opinions of the majority of participants, based on past conversations. By assigning different weights to these criteria for each agent, the framework can represent diverse intrinsic motivations. For instance, an agent with a higher weight on Novelty will prioritize curiosity and innovation, whereas an agent with a higher weight on Consensus will emphasize collaboration and agreement.
During the introspection process, the generated idea are critically examined to identify any technical ambiguities or contradictions using a language model. The internal state module then adjusts the idea to address these issues, refining it to improve clarity and coherence.
The internal state module not only reflects on the agent’s own thoughts about technological domains but also evaluates the opinions expressed by other participants. Specifically, based on prior conversations and the agent’s current thoughts, it identifies aspects of each participant’s opinions that the agent agrees with or finds interesting, as well as those that the agent considers contradictory or lacking.
As a result, the agent’s current thoughts are continuously updated, incorporating refined ideas and reflections on others. The format of the current thoughts is given in Appendix -A. During the dialogue, the agent makes statements based on its current thoughts, shaped through these processes, as well as on memories retrieved from the memory module. These current thoughts are also used when reflecting on new information or statements acquired or heard during the dialogue. Even if agents are initially assigned homogeneous roles and personality settings, their thoughts dynamically evolve in distinct ways through these processes.
II-B Dialogue scheme for Analogy-driven innovation
Innovation takes various forms, such as technology-driven innovation, market-driven innovation, open, or closed innovation. This paper focuses on Design by Analogy (DbA), an approach to innovation that has been extensively studied across disciplines, including management science and cognitive science ([holyoak1994, gassmann2008opening, linsey2008modality, fu2015design, fu2013meaning, mcadams2002quantitative, song2022design, song2019design, hope2017accelerating, jeong2014creating, moreno2014fundamental, murphy2014function, linsey2008increasing, goel1997design, huhns1988argo, Melnychuk2020]). Many groundbreaking innovations, such as the application of racing car technologies to running shoes and various forms of biomimicry, employ cross-domain analogies. DbA is a method that addresses problems in a target domain by extracting and transferring design solutions from other domains (source domains). The central focus of this approach lies in identifying similarities and differences. In particular, it is widely recognized that superior innovations abstract away from surface-level or attribute-based similarities, instead relying on functional and relational similarities for analogy ([holyoak1994, linsey2008modality, fu2015design, fu2013meaning, mcadams2002quantitative, bhatta1996design]).
Building on these interdisciplinary insights, this study proposes a dialogue scheme that leverages collective reasoning among agents to replicate the process of DbA, extracting functional similarities and mechanical differences from technical documents and transferring solutions across domains.
Specifically, this scheme facilitates the following stages in sequence:
Preparation: Reviewing technical documents from different domains. Each agent is provided with multiple technical documents belonging to the target domain and source domains and is tasked with reviewing them.
Phase A: Extracting functional similarities. Through multiple rounds of discussion, the agents identify functional similarities between the target domain and the source domains. They also determine which of these similarities are more central to the problem at hand.
Phase B: Extracting mechanical differences. Building on the functional similarities identified in Phase A, the agents examine the mechanisms that generate these shared functions in the target and source domains. Through several rounds of discussion, they identify the differences in these mechanisms and evaluate which of these differences are most significant.
Phase C: Transferring solutions. In this phase, the agents explore how the mechanical differences identified in Phase B can be applied to address the challenges in the target domain. Through multiple rounds of discussion, they determine what aspects of the source domains’ mechanisms need to be modified and incorporated into the target domain, as well as what features of the target domain’s existing functions or mechanisms should be retained or discarded.
Phase D: Identifying new challenges. As the agents deliberate on the solution transfer discussed in Phase C, they also assess whether any unforeseen challenges might arise during the application of these solutions.
Phase E: Identifying new opportunities. In contrast to Phase D, this phase focuses on exploring whether the solution transfer creates new value beyond solving the existing challenges in the target domain.
A sample of the dialogue prompts designed for facilitating this process is provided in Appendix -B.
II-C Organizational Structure
The organizational structure of the agent team is defined as a directed graph, with agents as vertices and the directions of communications or instructions as edges. Each agent can only communicate with adjacent agents and issue instructions in the direction of the edges. By utilizing this directed graph structure, the framework can represent various organizational forms, ranging from flat to hierarchical structures, and evaluate their impact on technological discovery.
III Experiment
In the experiment, we use Dyson’s invention of the bladeless household fan, Dyson Air Multiplier, as an example of innovation and evaluate the extent to which the proposed framework can replicate the key elements of this invention’s ideas. The evaluation is conducted in a sandbox environment using fictional technical documents.
III-A Case: Dyson Air Multiplier
In household fans, it is common to cover the spinning blades with a grid-like guard to prevent accidental contact. However, such guards not only reduce the efficiency and consistency of airflow but also make the device less convenient to carry and clean. Dyson Air Multiplier addressed these challenges in household fans by applying the principles of fluid dynamics used in jet engines to household fans and achieving a transition from a physical blade-based airflow mechanism to a fluid dynamics-based airflow mechanism utilizing pressure differential.
The Dyson Air Multiplier is often referred to as a ”bladeless fan,” but this label is somewhat misleading ([hallgeisler2022]). In reality, the device uses blades enclosed within the base, which are powered by an efficient brushless motor to draw in air. Atop the base, a ring-shaped airfoil with a diameter of approximately 30 cm is mounted. Air drawn into the device is expelled through the narrow slits of a jet nozzle on the inner surface of the ring. During this process, the fan pulls in air from behind the ring, effectively ”multiplying” the airflow ([mansson2014flow]). Specifically, the air drawn from the base is accelerated through the jet nozzle, creating a pressure gradient that induces surrounding air to be pulled in (inducement). Additionally, the high-speed expelled air entrains the surrounding air, generating a consistent airflow directed forward from the ring (entrainment).
Dyson introduced the first generation of the Dyson Air Multiplier to the market in 2009. However, the first-generation device faced challenges with noise levels. To address this issue, Dyson invested over \$60 million in further research and development, culminating in the release of the second generation in 2014. The second-generation model incorporated Helmholtz cavities into the base, effectively resolving the turbulence issues that were a primary source of noise ([guardian2014, hallgeisler2022]).
III-B Experimental Method
In the experiment, dialogues among generative agents are conducted with household fans as the target domain and industrial ejectors as the source domain. An ejector is an industrial device used for transporting fluids (liquids or gases). It is utilized not only for water supply, exhaust, and smoke extraction but also for neutralization, dilution, and cooling through the mixing of gases or liquids, as well as for creating vacuum conditions as a vacuum pump.
Although the application domains are entirely different, household fans and industrial ejectors share a functional similarity in the movement of fluids. However, the mechanisms that enable this function are entirely distinct. In a typical ejector, a nozzle draws high-speed driving fluid into the body as the driving source. As the high-speed driving fluid reduces the pressure inside the body, creating negative pressure, the surrounding low-speed external fluid is pulled into the body. The mixed fluid of the driving and external fluids is then discharged outside the body through a diffuser. This allows the ejector to move external fluid without requiring any external power source.
In the practical stage of the proposed framework, actual technical documents, such as patent filings, are expected to be used. However, in the sandbox environment presented in this paper, fictional technical documents based on real-world technical information are prepared. These fictional documents, like actual patent documents, include miscellaneous information that does not necessarily lead to the target invention. For instance, in the case of industrial ejectors, the document contains information about a wide range of industrial applications far removed from household use, details about unrelated components, and data on various types of irrelevant fluids and materials. Similarly, the technical documents for household fans include various unrelated information. Agents must extract fragments of information that lead to a ”solution” from this broad and miscellaneous data and connect those fragments using their independent reasoning. In this particular experiment, the agents must reinterpret the concept of driving fluid from the perspective of a household fan and devise a method to generate it within the constraints of household technology and energy environments. However, the agents are prohibited, by prompts, from using pre-trained knowledge about Dyson products or any information available on the internet. Instead, they are required to base their statements and reflections strictly on logical conclusions or associations derived from their prior discussions and reflections.
III-C Evaluations
The evaluation is conducted for the solutions derived by the agents from the dialogue using a point-based system based on the nine criteria listed in Table I (maximum of 8 points). These criteria are developed from the perspective of whether the transfer of solutions from industrial ejectors to household fans comprehensively and coherently covers the technological elements corresponding to the first-generation Dyson Air Multiplier. Since the elements of the second generation may require computational fluid dynamics simulations for the specific shape of the fan to address the turbulence issues, they were excluded from the evaluation in this experiment, which only utilizes language models.
Criterion Description
1 If no direct mention of blades is made, is a realistic and safe device capable of generating driving fluid within the size and price range of household appliances proposed? (2 points)
2 Is retaining blades proposed, and are they identified as playing a role in generating driving fluid or negative pressure? (1 point)
3 Is retaining blades proposed, and is enclosing them inside the main body suggested? (1 point)
4 If no direct mention of a motor is made, is retaining blades proposed, or is a device as described in criterion 1 conventionally paired with a motor proposed? (1 point) Additionally, is the use of household electricity as the motor’s power source mentioned? (1 point)
5 Is retaining the motor proposed, and is it identified as playing a role in generating driving fluid or negative pressure? (1 point)
6 Are advanced proposals made regarding a motor to improve its efficiency or noise reduction, such as the use of a brushless motor? (1 point)
7 Is the introduction of a nozzle structure proposed? (1 point)
8 Other than the introduction of a nozzle structure, are any proposals made to optimize internal channels for improving the efficiency or noise reduction of the driving fluid? (Up to 2 points, depending on the number and specificity of the proposals)
9 Are relevant principles of fluid dynamics explicitly explained? (1 point)
TABLE I: Evaluation criteria
To evaluate the performance of the proposed framework, the experimental results are compared across eight models shown in Table II. The eight models are differentiated based on three perspectives: (1) the number of agents, (2) the presence of internal states, and (3) the heterogeneity of intrinsic motivations. The first perspective examines the impact of increasing the number of agents participating in the dialogue. When multiple agents are involved, the evaluation is conducted based on the majority opinions. The second perspective assesses the role of the internal state module’s functionality. The third perspective evaluates the effect of whether the weights assigned to the three intrinsic motivations (novelty, importance, and consensus) are homogeneous or heterogeneous among agents. In homogeneous models (Models D and G), all agents equally consider the three motivations. In heterogeneous models (Models E and H), agents are configured with different emphases on intrinsic motivations: some prioritize novelty while de-emphasizing consensus, others prioritize consensus while de-emphasizing novelty, and some consider all three motivations equally.
Model Number of agents Internal state intrinsic motivation
A: 1 no-internal agent 1 No n.a.
B: 1 with-internal agent 1 Yes n.a.
C: 3 no-internal agents 3 No n.a.
D: 3 with-internal homogeneous agents 3 Yes homogeneous
E: 3 with-internal heterogeneous agents 3 Yes heterogeneous
F: 5 no-internal agents 5 No n.a.
G: 5 with-internal homogeneous agents 5 Yes homogeneous
H: 5 with-internal heterogeneous agents 5 Yes heterogeneous
TABLE II: Eight models for experiment
The experiment is conducted 10 times for each of the eight models, using a common set of 10 seeds. The language model used to power the framework is OpenAI’s GPT-4o. To ensure uniform experimental conditions, word limits are imposed on the agents’ statements in each round. When expressing their own opinions, agents are limited to 170 words, while statements regarding the opinions of other agents are restricted to 50 words. In this preliminary experiment, a flat organizational structure was adopted, where all agents could communicate with all other agents. Other organizational structures, such as hierarchical models, are left for future research. In the evaluation, the ratings are done by carefully examining each agent’s statement on possible solutions, one by one, through human assessment.
III-D Results
As shown in Table III, in all cases with 1, 3, and 5 agents, the models with internal states outperformed those without on average. This indicates that, through refinement by the internal state module, more coherent and comprehensive ideas are constructed and shared among agents. For example, comparing the two models with one agent, in the model without an internal state (Model A), the proposal of retaining blades or incorporating an appropriate blade-driven device as a mechanism to generate a driving fluid or negative pressure (Criteria 1 or 2) was made only 3 times out of 10. In contrast, in the model with an internal state (Model B), such mechanisms were proposed 9 times out of 10. This indicates, in the former model, the agent’s reflection did not address how to deal with the external driving fluid required in incorporating ejector-like technology into household fan, whereas in the latter model, the concept of external driving fluid was successfully reframed within the context of household fans as the intake of surrounding air by the fan blades. Furthermore, in the model without an internal state, aside from the introduction of nozzle structures, the proposal for optimizing internal channels to improve the efficiency and reduce the noise of the driving fluid (Criterion 8) was made only once out of 10 times. In contrast, in the model with an internal state, such proposals were made 7 times out of 10, ranging from the redesign of the fan housing to the introduction of guide vanes and the use of Computational Fluid Dynamics (CFD) simulations. Below is an example of part of an agent’s statement in the single-agent model with an internal state (Model B).
To address challenges in household fans, we could integrate concepts from ejector mechanisms while selectively retaining and discarding traditional fan elements. ... Rather than relying solely on large rotating blades, we might consider smaller impellers driven by the existing motor, which would introduce directed air streams. ... Retaining a compact motor system that uses lightweight, efficient turbines is crucial, as it allows for consistent airflow without excessive bulk. ... By strategically designing air inlets and channels, we can naturally guide air currents, even without a high-velocity driving fluid. Prioritizing safety, all moving parts should be enclosed within the casing, utilizing protective materials to eliminate the risk of contact with blades. ... Utilize computational fluid dynamics (CFD) to refine blade shapes and angles, enhancing air movement and improving efficiency. ... Incorporate guide vanes that naturally direct airflow, minimizing reliance on high-velocity inputs. This approach ensures smoother air pressure control and efficient cooling without adding mechanical complexity. ...
In contrast,the interpretation of the impact of the number of agents is not straightforward. On one hand, the effect is generally positive for models without an internal state. First, the average performance improved significantly when the number of agents increased from one (Model A) to multiple (Models C and F). Second, when comparing the three-agent model (Model C) to the five-agent model (Model F), increasing the number of agents significantly reduced the variance, although the average scores remained unchanged. The larger variance in scores with three agents may partly stem from the stricter majority requirement for consensus, where the opinion of a single agent can heavily influence the results.
On the other hand, when the agents have an internal state, the single agent model (Model B) outperformed almost all multiple-agent models both with and without an internal state (Models C, D, E, F, and G) except for Model H. It also recorded the smallest variance across all models. This highlights the strong capability of the internal state module in deriving comprehensive and coherent ideas, even when acting alone. However, its superior performance compared to some multi-agent models with internal states may also be attributed to the stricter majority requirement. In particular, the larger variance observed in the scores with three agents suggests this.
The highest average score among all models was recorded by the case with five heterogeneous agents with an internal state (Model H). Increasing the number of agents from 3 to 5 slightly relaxed the consensus requirement, allowing a broader range of opinions to emerge from more diverse agents, which were then refined and shared through collective reasoning. The positive effects of the number and diversity of agents are consistent with the results of some prior studies on LLM-MAAS ([du2023improvingfactualityreasoninglanguage, chan2023chatevalbetterllmbasedevaluators]). Below is an example of part of agents’ statement in the multiple-agent model with an internal state (Model H).
Agent_1: ... First, retaining the electric motor is crucial, but shifting the focus from blades to a ducted design with specialized nozzles could minimize safety risks and maintenance. ... This means adopting a circular or linear duct system to channel airflow, creating efficient fluid movement pathways. The key is to enhance the fan’s ability to draw in and expel air smoothly using controlled high-speed airflow, perhaps with adjustable nozzles or diffusers that modify air velocity and direction. ...
Agent_2: ... However, traditional metal or plastic grids could be phased out if the airflow generation can be enclosed entirely within the fan’s structure.
Agent_3  ... The most critical adaptation would be incorporating the ejector’s pressure-driven flow to create a bladeless fan. This could be achieved using a small compressor or blower to generate a high-pressure air source, channeling it through a nozzle to produce a coherent airflow. ...
Agent_5: ... Retaining the motor to power a compressor or blower generating a directed airflow stream aligns with fluid dynamics principles, enhancing efficiency. ...
Agent_1: ... Noise minimization can be achieved by including sound-dampening materials and CFD-aided design to manage high-speed airflow noise impacts. ...
The agents’ dialogue above proposed several ideas, including eliminating the existing grid guard and conducting airflow generation within an enclosed environment inside the fan, adopting a ducted design to improve airflow efficiency, utilizing a motor-driven blower to generate high-speed driving fluid, and performing CFD simulations to reduce noise. These proposals demonstrate that the key ideas underlying the Dyson Air Multiplier have been successfully replicated.
Model Min Max Average Standard deviation
A: 1 no-internal agent 1 5 2.4 1.58
B: 1 with-internal agent 4 7 5.1 0.99
C: 3 no-internal agents 1 7 3.7 2.45
D: 3 with-internal homogeneous agents 2 7 4.4 2.07
E: 3 with-internal heterogeneous agents 2 7 4.0 2.05
F: 5 no-internal agents 1 6 3.7 1.83
G: 5 with-internal homogeneous agents 2 6 4.7 1.16
H: 5 with-internal heterogeneous agents 3 7 5.3 1.16
TABLE III: Experimental results
IV Conclusion
This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation. To achieve this, it proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of analogy-driven innovation. The framework’s functionality is evaluated using Dyson’s invention of the bladeless household fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through a set of fictional technical documents.
The experimental results demonstrate that models with internal states significantly outperformed those without, achieving higher average scores and lower variance. Notably, the model with five heterogeneous agents equipped with internal states successfully replicated the key ideas underlying the Dyson Air Multiplier. This indicates that the internal state module enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
-A Format of user’s current thought
a. Functional Similarlities
The functional similarities between the target domain and the ejector:
- In both domains, ...
The points in agent_2’s opinion on the topic of the functional similarities that I agree with or feel interesting are as follows:
The points in agent_2’s opinion on the topic of the functional similarities that contain contradictions or missing information are as follows:
b. Mechanical Differences
How the mechanisms in the ejector that produce the common functionalities differ from those in the target domain:
c. Solutions
How the functionalities or mechanisms of the ejector and the target domain should be modified or retained to solve the challenges of the target domain:
d. New Challenges
The potential new challenges arising from applying the ejector’s mechanisms:
e. New Opportunities
The additional opportunities generated by applying the ejector’s mechanisms:
-B Sample prompts for dialogue
Phase A: Functional similarities
Hello. The field of the product we are developing is referred to as the ’target domain,’ and the field of reference products for development is referred to as the ’source domain.’ The target domain for this session is household fan. The source domain is ejector. Read the following documents about the target domain and the source domain, and provide your thoughts on the core functional similarities between the two domains, starting with what you believe to be the most essential similarities. However, do not discuss the differences between the two domains, the challenges within each domain, their possible solutions yet. Please respond within 170 words.
Phase B: Mechanical differences
Thank you. Next, please share your thoughts on how the mechanisms in the source domain that produce the common functionalities differ from those in the target domain, starting with the most significant differences. However, do not discuss the challenges within each domain or their solutions yet. Please respond within 170 words.
Phase C: Solutions
Thank you. Now, let’s consider how applying the discussed differences in mechanisms to the target domain could help solve its challenges. Please share your thoughts on what aspects of the source domain’s mechanisms should be adapted and how they should be integrated into the target domain to address the challenges in the target domain. Additionally, explain what functionalities or mechanisms of the target domain should be retained and what should be discarded, in order of importance, starting with the most critical points. Please respond within 170 words.
Phase D: New challenges
Thank you. Now, let’s consider whether applying the source domain’s mechanisms to the target domain in the manner discussed could lead to any unforeseen significant challenges. If you believe such challenges might arise, please share your thoughts, listing them in order of severity. If you do not believe any major challenges would arise, please state that as your answer. Please respond within 170 words.
Phase E: New opportunities
Thank you. Now, let’s consider the opposite perspective. Could applying the source domain’s mechanisms to the target domain create additional opportunities beyond merely solving the target domain’s existing challenges? If you believe such new opportunities exist, please share your thoughts, listing them in order of importance. If you do not believe any additional opportunities would be generated, please state that as your answer. Please respond within 170 words.
[…]

### v3

@@
-GAI_arXiv.bib
+
 Abstract
 This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation.
-To achieve this, it proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of innovation.
+To achieve this, it proposes GAI (Generative Agents for Innovation), a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of innovation.
 The core of the GAI framework lies in an architecture that dynamically processes the internal states of agents and a dialogue scheme specifically tailored to facilitate analogy-driven innovation.
-The framework’s functionality is evaluated using Dyson’s invention of the bladeless fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through a set of fictional technical documents.
+The framework’s functionality is evaluated using Dyson’s invention of the bladeless fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through agents’ dialogue using a set of fictional technical documents.
 The experimental results demonstrate that models with internal states significantly outperformed those without, achieving higher average scores and lower variance.
 Notably, the model with five heterogeneous agents equipped with internal states successfully replicated the key ideas underlying the Dyson’s invention.
-This indicates that the internal state enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
-Keywords: generative AI, Large language models, Multi-Agent Systems, Innovation, Patent
-I Introduction
-Generative agents, including LLM agents, can engage in reflections for decision-making or analysis by breaking down complex tasks and conducting systematic reasoning.
-LLM-based Multi-agent System (LLM-MAS) extends this capability by adding collective reasoning functions, with applications already emerging in fields such as software development and social simulations ([chan2023chatevalbetterllmbasedevaluators, guo2024largelanguagemodelbased, hong2024metagptmetaprogrammingmultiagent, park2023generativeagentsinteractivesimulacra, qian2024chatdevcommunicativeagentssoftware, sun2024llmbasedmultiagentreinforcementlearning, Xiong_2023, zhou2024sotopiainteractiveevaluationsocial]).
-One particularly noteworthy area is the application of LLM-MAS to scientific discussions.
-In some studies, generative agents have been used to address tasks including mathematical problems and medical debates, with interesting findings such as agents correcting their mistakes through dialogue ([tang2024medagentslargelanguagemodels, du2023improvingfactualityreasoninglanguage]).
-Applying these capabilities of generative agents to discussions based on technical documents, such as patents, could potentially replicate the collective interactions that human teams engage in during the innovation process.
-Efforts to analyze the innovation process have historically included observational experiments in actual corporate environments, but these have faced significant limitations due to low reproducibility and difficulties in manipulating experimental conditions.
-As a result, the effects of variables such as organizational structure, communication methods, and degree of diversity on the innovation process remain not fully understood.
-If LLM-MAS can replicate the innovation process, it would enable repeated simulations with adjustable conditions under high reproducibility.
-Furthermore, it is conceivable that artificial intelligence technologies, including LLMs, will increasingly be applied in real-world corporate innovation settings, creating a new domain of practice that could be described as “AI-driven innovation.” If LLM-MAS can simulate the innovation process, it would also be possible to pioneer this new domain by integrating human teams with generative agent teams, demonstrating new possibilities for AI-human collaboration in driving innovation.
+This indicates that the framework enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
+{IEEEkeywords}
+Generative AI, Innovation, Large language models, Multi-Agent Systems, Patents
+1 INTRODUCTION
+Generative agents, including LLM agents, can perform various complex tasks by breaking them down, conducting systematic reasoning, leveraging external resources, learning from experiences, and planning for future.
+LLM-based Multi-agent System (LLM-MAS) extends this capability by incorporating collective reasoning functions, with applications already emerging in fields such as software development, urban planning, scientific debates, and social simulations ([1, 7, 8, 9, 16, 12, 19, 32, 35, 41, 43, 44, 45, 46, 48, 49, 53, 65, 67, 68, 73, 76, 80, 86, 87, 88, 90, 91, 96, 95, 98, 103]).
+Applying these capabilities of generative agents to discussions based on technical documents, such as patents, could potentially replicate the collective interactions that human teams engage in during innovation processes.
+Efforts to analyze innovation processes have historically included observational experiments in actual corporate environments, but these have faced significant limitations due to low reproducibility and difficulties in manipulating experimental conditions.
+As a result, the effects of conditions such as diversity in team member, communication methods, and organizational structure on innovation processes remain not fully understood.
+If LLM-MAS can replicate the innovation processes of human teams, it would enable repeated simulations under adjustable conditions with high reproducibility.
+Furthermore, artificial intelligence technologies, including LLMs, will increasingly be applied in real-world corporate innovation settings ([2, 4, 5, 6, 13, 15, 22, 23, 29, 30, 31, 57, 69, 71, 74, 75, 81, 82, 94]).
+If LLM-MAS can simulate innovation processes, it would also be possible to pioneer this new domain by integrating generative agent teams with human teams, demonstrating new possibilities for AI-human collaboration in driving innovation.
 The objective of this study is to examine whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation.
-To achieve this, the study specifically focuses on analogy-driven innovation and proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate this process.
-The core of the GAI framework lies in an architecture that dynamically processes the internal states of agents and a dialogue scheme specifically designed to facilitate analogy-driven innovation.
-In the former, agents generate and refine ideas based on intrinsic motivation and critical introspection, leveraging previously reviewed documents and interactions with others to form their current thoughts, which serve as the basis for their contributions to the dialogue.
-In the latter, the scheme incorporates insights from management science and cognitive science to replicate the process of identifying functional similarities and mechanical differences, enabling the transfer of solutions across diverse technical domains.
-To evaluate the functionality of the framework, this study uses Dyson’s invention of the bladeless household fan as a case study.
-It examines the extent to which analogy-driven innovations can be replicated by the proposed framework using a set of fictional technical documents.
-II Proposed framework
-II-A Agent architecture
-The architecture of each agent consists of two primary modules driven by an LLM: a memory module and an internal state module (Figure 1).
-The memory module records the contents of reviewed technical documents, past conversations with others, and the agent’s prior reflections, all timestamped for reference.
-These records are then organized into a structured memory, which is input into the internal state module.
-The internal state module processes newly acquired technical information and statements from others during the dialogue, drawing on both recent thoughts and memories retrieved from the memory module.
-Based on this, the internal state module forms and updates the agent’s current thoughts, which serve as the foundation for its future reflections and statements.
+To achieve this, the study focuses on analogy-driven innovation and proposes GAI (Generative Agents for Innovation), a new LLM-empowered framework designed for reflection and interaction among generative agents to replicate this process.
+The core of the GAI framework lies in an agent’s architecture that dynamically processes its internal state, and a dialogue scheme specifically designed to facilitate analogy-driven innovation.
+In the former, each agent generates and refines ideas based on intrinsic motivation and critical introspection, leveraging previously reviewed documents and interactions with others.
+This process forms and updates the current thought of the agent, allowing the agent’s contributions to the dialogue to dynamically self-evolve in a coherent way over time.
+In the latter, the scheme incorporates insights from management and cognitive science to replicate the process of identifying functional similarities and mechanical differences, enabling the transfer of solutions across various technological domains.
+To evaluate the functionality of the framework, this study performs two experiments using a set of fictional technical documents.
+The one is to take Dyson’s invention of the bladeless fan as a case study and examine the extent to which the innovation can be replicated using household fans as the target domain and industrial ejectors as the source domain.
+The other is to apply the framework to ten different source domains to examine its applicability.
+2 PROPOSED FRAMEWORK
+2.1 AGENT ARCHITECTURE
+The architecture of an agent consists of two primary LLM-driven modules: a memory module and an internal state module (Figure 1).
+The memory module records a) contents of reviewed technical documents, b) past conversations with others, and c) the agent’s prior reflections, all timestamped for reference.
+These records are then organized into structured memory, which is input into the internal state module.
+The internal state module processes newly acquired technical information and statements from others during the dialogue and facilitates innovative and coherent thinking of an agent.
+Specifically, taking in external inputs from technical documents and other agents, as well as internal inputs from the agent’s prior reflections and its latest thoughts, the module continuously updates the agent’s current thoughts, which is a collection of texts structured according to the dialogue scheme for analogy-driven innovation.
+Current thoughts recurrently serve as the foundation for future reflections of the agent and function as the basis for the agent’s statement in a dialogue, providing it with strong context awareness and coherency.
 Figure 1: Agent architecture and organizational structure
 When forming and updating current thoughts, the internal state module undergoes two processes: idea generation and introspection.
-During the idea generation process, the module generates multiple ideas for reflections on a technological domain based on the memories received from the memory module and the agent’s thoughts at that time.
-Subsequently, the internal state module evaluates these ideas on a scale of 1 to 10 for each of three criteria—Novelty, Importance, and Consensus—using a language model and selects ideas that maximize an intrinsic reward calculated as a weighted sum of these criteria:
+During the idea generation process, the module is prompted to generate multiple ideas distinct with each other for reflections on a technological domain based on the memories retrieved from the memory module and the agent’s current thoughts at that time.
+Subsequently, the module initiates another LLM instance and prompts it to evaluate these ideas by an integer score on a scale of 1 to 10 for each of three criteria:
+  • Novelty: the degree to which the generated idea introduces new content that has not been present in prior conversations and reflections
+  • Importance: the degree to which the generated idea is significant in light of previous discussions and reflections
+  • Consensus: the degree to which the generated idea aligns with the opinions of the majority of participants, based on past conversations
+The module selects the idea that maximizes an intrinsic reward calculated as a weighted sum of these criteria:
 $\begin{split}\max_{x_{ij}}r_{ij}&\equiv\alpha_{i}\,\text{Novelty}(x_{ij})+\beta_{i}\,\text{Importance}(x_{ij})\\ &+\gamma_{i}\,\text{Consensus}(x_{ij}),\end{split}$ (1)
-where $r_{ij}$ represents the intrinsic reward of agent $i$, $x_{ij}$ is its idea on object $j$, and $\alpha_{i}$, $\beta_{i}$, and $\gamma_{i}$ respectively represent the weights for the 1-10 evaluation on each criterion, which are denoted by Novelty$(x_{ij})$, Importance$(x_{ij})$, and Consensus$(x_{ij})$ in Equation (1).
-The Novelty criterion assesses the degree to which an idea introduces new content that has not been present in prior conversations and reflections.
-The Importance criterion evaluates how significant the idea is in light of previous discussions and reflections.
-The Consensus criterion measures the extent to which the idea aligns with the opinions of the majority of participants, based on past conversations.
-By assigning different weights to these criteria for each agent, the framework can represent diverse intrinsic motivations.
-For instance, an agent with a higher weight on Novelty will prioritize curiosity and innovation, whereas an agent with a higher weight on Consensus will emphasize collaboration and agreement.
-During the introspection process, the generated idea are critically examined to identify any technical ambiguities or contradictions using a language model.
-The internal state module then adjusts the idea to address these issues, refining it to improve clarity and coherence.
+[…]
+During the introspection process, the selected idea is critically examined to identify any technical ambiguities or contradictions.
+Specifically, the module sets up another LLM instance as an internal critic and prompts it to examine if there are any important points in the selected idea that are unclear from the perspective of technology, and if there are any contradictions from the perspective of technology.
+The internal state module is then prompted to update the idea by including responses that accurately address the questions and points of contradiction raised through the introspection, refining the idea to improve clarity and coherence.
 The internal state module not only reflects on the agent’s own thoughts about technological domains but also evaluates the opinions expressed by other participants.
-Specifically, based on prior conversations and the agent’s current thoughts, it identifies aspects of each participant’s opinions that the agent agrees with or finds interesting, as well as those that the agent considers contradictory or lacking.
-As a result, the agent’s current thoughts are continuously updated, incorporating refined ideas and reflections on others.
-The format of the current thoughts is given in Appendix -A.
+Specifically, based on prior conversations and the agent’s current thoughts, the module is prompted to identify aspects of each participant’s opinions that the agent agrees with or finds interesting, as well as those that it considers contradictory or lacking.
+These reflections on others are then reflected in the agent’s current thoughts together with its own ideas.
+Through these steps, the agent’s current thoughts are continuously updated, incorporating refined ideas and reflections on domains and others.
 During the dialogue, the agent makes statements based on its current thoughts, shaped through these processes, as well as on memories retrieved from the memory module.
-These current thoughts are also used when reflecting on new information or statements acquired or heard during the dialogue.
-Even if agents are initially assigned homogeneous roles and personality settings, their thoughts dynamically evolve in distinct ways through these processes.
-II-B Dialogue scheme for Analogy-driven innovation
-Innovation takes various forms, such as technology-driven innovation, market-driven innovation, open, or closed innovation.
-This paper focuses on Design by Analogy (DbA), an approach to innovation that has been extensively studied across disciplines, including management science and cognitive science ([holyoak1994, gassmann2008opening, linsey2008modality, fu2015design, fu2013meaning, mcadams2002quantitative, song2022design, song2019design, hope2017accelerating, jeong2014creating, moreno2014fundamental, murphy2014function, linsey2008increasing, goel1997design, huhns1988argo, Melnychuk2020]).
+Even if agents are initially assigned homogeneous roles and personality, their thoughts dynamically evolve in distinct ways through these processes.
+2.2 DIALOGUE SCHEME FOR ANALOGY-DRIVEN INNOVATION
+Innovation takes various forms, such as technology-driven innovation, market-driven innovation, open or closed innovation.
+This paper focuses on Design by Analogy (DbA), an approach to innovation that has been extensively studied across disciplines, including management science and cognitive science ([17, 18, 20, 25, 42, 33, 37, 39, 50, 51, 58, 59, 61, 63, 77, 78]).
 Many groundbreaking innovations, such as the application of racing car technologies to running shoes and various forms of biomimicry, employ cross-domain analogies.
-DbA is a method that addresses problems in a target domain by extracting and transferring design solutions from other domains (source domains).
+DbA addresses problems in a target domain by extracting and transferring design solutions from other domains (source domains).
 The central focus of this approach lies in identifying similarities and differences.
-In particular, it is widely recognized that superior innovations abstract away from surface-level or attribute-based similarities, instead relying on functional and relational similarities for analogy ([holyoak1994, linsey2008modality, fu2015design, fu2013meaning, mcadams2002quantitative, bhatta1996design]).
+In particular, it is widely recognized that superior innovations abstract away from surface-level or attribute-based similarities, instead relying on functional and relational similarities for analogy ([3, 18, 17, 42, 51, 58]).
 Building on these interdisciplinary insights, this study proposes a dialogue scheme that leverages collective reasoning among agents to replicate the process of DbA, extracting functional similarities and mechanical differences from technical documents and transferring solutions across domains.
-Specifically, this scheme facilitates the following stages in sequence:
-Preparation: Reviewing technical documents from different domains.
-Each agent is provided with multiple technical documents belonging to the target domain and source domains and is tasked with reviewing them.
-Phase A: Extracting functional similarities.
-Through multiple rounds of discussion, the agents identify functional similarities between the target domain and the source domains.
-They also determine which of these similarities are more central to the problem at hand.
-Phase B: Extracting mechanical differences.
-Building on the functional similarities identified in Phase A, the agents examine the mechanisms that generate these shared functions in the target and source domains.
-Through several rounds of discussion, they identify the differences in these mechanisms and evaluate which of these differences are most significant.
-Phase C: Transferring solutions.
-In this phase, the agents explore how the mechanical differences identified in Phase B can be applied to address the challenges in the target domain.
-Through multiple rounds of discussion, they determine what aspects of the source domains’ mechanisms need to be modified and incorporated into the target domain, as well as what features of the target domain’s existing functions or mechanisms should be retained or discarded.
-Phase D: Identifying new challenges.
-As the agents deliberate on the solution transfer discussed in Phase C, they also assess whether any unforeseen challenges might arise during the application of these solutions.
-Phase E: Identifying new opportunities.
-In contrast to Phase D, this phase focuses on exploring whether the solution transfer creates new value beyond solving the existing challenges in the target domain.
-A sample of the dialogue prompts designed for facilitating this process is provided in Appendix -B.
-II-C Organizational Structure
+Specifically, this scheme facilitates the procedure shown in Figure 2.
+Figure 2: Dialogue scheme for analogy-driven innovation
+2.3 ORGANIZATIONAL STRUCTURE
 The organizational structure of the agent team is defined as a directed graph, with agents as vertices and the directions of communications or instructions as edges.
@@
 By utilizing this directed graph structure, the framework can represent various organizational forms, ranging from flat to hierarchical structures, and evaluate their impact on technological discovery.
-III Experiment
-In the experiment, we use Dyson’s invention of the bladeless household fan, Dyson Air Multiplier, as an example of innovation and evaluate the extent to which the proposed framework can replicate the key elements of this invention’s ideas.
-The evaluation is conducted in a sandbox environment using fictional technical documents.
-III-A Case: Dyson Air Multiplier
+3 EXPERIMENTS
+3.1 CASE: Dyson Air Multiplier
 In household fans, it is common to cover the spinning blades with a grid-like guard to prevent accidental contact.
 However, such guards not only reduce the efficiency and consistency of airflow but also make the device less convenient to carry and clean.
-Dyson Air Multiplier addressed these challenges in household fans by applying the principles of fluid dynamics used in jet engines to household fans and achieving a transition from a physical blade-based airflow mechanism to a fluid dynamics-based airflow mechanism utilizing pressure differential.
-The Dyson Air Multiplier is often referred to as a ”bladeless fan,” but this label is somewhat misleading ([hallgeisler2022]).
-In reality, the device uses blades enclosed within the base, which are powered by an efficient brushless motor to draw in air.
+Dyson Air Multiplier addressed these challenges applying the principles of fluid dynamics used in jet engines and achieving a transition from a physical-blade-based airflow mechanism to a fluid-dynamics-based airflow mechanism leveraging pressure differential.
+Air Multiplier is often referred to as a ”bladeless fan,” but this is somewhat misleading ([28]).
+In reality, it uses blades enclosed within the base, powered by an efficient brushless motor to draw in air.
 Atop the base, a ring-shaped airfoil with a diameter of approximately 30 cm is mounted.
-Air drawn into the device is expelled through the narrow slits of a jet nozzle on the inner surface of the ring.
-During this process, the fan pulls in air from behind the ring, effectively ”multiplying” the airflow ([mansson2014flow]).
+Air drawn into the device is expelled through narrow slits of the jet nozzle on the inner surface of the ring.
+During this process, the fan pulls in air from behind the ring, effectively ”multiplying” the airflow ([56]).
 Specifically, the air drawn from the base is accelerated through the jet nozzle, creating a pressure gradient that induces surrounding air to be pulled in (inducement).
 Additionally, the high-speed expelled air entrains the surrounding air, generating a consistent airflow directed forward from the ring (entrainment).
-Dyson introduced the first generation of the Dyson Air Multiplier to the market in 2009.
-However, the first-generation device faced challenges with noise levels.
-To address this issue, Dyson invested over \$60 million in further research and development, culminating in the release of the second generation in 2014.
-The second-generation model incorporated Helmholtz cavities into the base, effectively resolving the turbulence issues that were a primary source of noise ([guardian2014, hallgeisler2022]).
-III-B Experimental Method
-In the experiment, dialogues among generative agents are conducted with household fans as the target domain and industrial ejectors as the source domain.
-An ejector is an industrial device used for transporting fluids (liquids or gases).
-It is utilized not only for water supply, exhaust, and smoke extraction but also for neutralization, dilution, and cooling through the mixing of gases or liquids, as well as for creating vacuum conditions as a vacuum pump.
-Although the application domains are entirely different, household fans and industrial ejectors share a functional similarity in the movement of fluids.
+Dyson introduced the first generation of Air Multiplier to the market in 2009.
+However, it faced challenges with noise levels.
+To address this issue, Dyson invested over \$60 million in further R&D, culminating in the release of the second generation in 2014, which incorporated Helmholtz cavities into the base, effectively resolving the turbulence issues that were a primary source of noise ([26, 28]).
+3.2 EXPERIMENT 1
+In the experiment, dialogues are conducted with household fans as the target domain and industrial ejectors as the source domain.
+Ejectors are industrial devices used for transporting fluids (liquids or gases).
+They are utilized not only for water supply, exhaust, and smoke extraction but also for neutralization, dilution, and cooling through the mixing of gases or liquids, as well as for creating vacuum conditions as a vacuum pump.
+Although the applications are entirely different, the tow domains share a functional similarity, the movement of fluids.
 However, the mechanisms that enable this function are entirely distinct.
-In a typical ejector, a nozzle draws high-speed driving fluid into the body as the driving source.
+In a typical ejector, a nozzle draws high-speed driving fluid into the body as a driving source.
 As the high-speed driving fluid reduces the pressure inside the body, creating negative pressure, the surrounding low-speed external fluid is pulled into the body.
 The mixed fluid of the driving and external fluids is then discharged outside the body through a diffuser.
-This allows the ejector to move external fluid without requiring any external power source.
+This allows the ejector to move external fluid without any external power sources.
 In the practical stage of the proposed framework, actual technical documents, such as patent filings, are expected to be used.
-However, in the sandbox environment presented in this paper, fictional technical documents based on real-world technical information are prepared.
+However, in the experimental environment presented in this paper, fictional technical documents based on real-world technical information are prepared, primarily considering budgetary constraints.
 These fictional documents, like actual patent documents, include miscellaneous information that does not necessarily lead to the target invention.
-For instance, in the case of industrial ejectors, the document contains information about a wide range of industrial applications far removed from household use, details about unrelated components, and data on various types of irrelevant fluids and materials.
-Similarly, the technical documents for household fans include various unrelated information.
+For instance, the document for ejectors contains information about a wide range of industrial applications far removed from household use, details about unrelated components, and data on various types of irrelevant fluids and materials.
+Similarly, the document for household fans include various unrelated information.
 Agents must extract fragments of information that lead to a ”solution” from this broad and miscellaneous data and connect those fragments using their independent reasoning.
-In this particular experiment, the agents must reinterpret the concept of driving fluid from the perspective of a household fan and devise a method to generate it within the constraints of household technology and energy environments.
+In this particular experiment, to reach the target innovation, agents are required to reinterpret the concept of driving fluid from the perspective of a household fan and devise a method to generate it within the constraints of household technology and energy environments.
 However, the agents are prohibited, by prompts, from using pre-trained knowledge about Dyson products or any information available on the internet.
 Instead, they are required to base their statements and reflections strictly on logical conclusions or associations derived from their prior discussions and reflections.
-III-C Evaluations
-The evaluation is conducted for the solutions derived by the agents from the dialogue using a point-based system based on the nine criteria listed in Table I (maximum of 8 points).
-These criteria are developed from the perspective of whether the transfer of solutions from industrial ejectors to household fans comprehensively and coherently covers the technological elements corresponding to the first-generation Dyson Air Multiplier.
-Since the elements of the second generation may require computational fluid dynamics simulations for the specific shape of the fan to address the turbulence issues, they were excluded from the evaluation in this experiment, which only utilizes language models.
-Criterion Description
+The evaluation is conducted for the solutions derived from the dialogue using a point-based system based on the nine criteria listed in Table 1 (maximum of 8 points).
+These criteria are developed from the perspective of whether the transfer of solutions comprehensively and coherently covers the technological elements corresponding to the first-generation Dyson Air Multiplier.
+Since the second generation requires computational fluid dynamics simulations for the specific shape of the fan to address the turbulence issues, they were excluded from the evaluation in this experiment, which only utilizes language models.
+Table 1: Evaluation criteria
+No. Description
 1 If no direct mention of blades is made, is a realistic and safe device capable of generating driving fluid within the size and price range of household appliances proposed? (2 points)
@@
 3 Is retaining blades proposed, and is enclosing them inside the main body suggested? (1 point)
-4 If no direct mention of a motor is made, is retaining blades proposed, or is a device as described in criterion 1 conventionally paired with a motor proposed?
+4 If no direct mention of a motor is made, is retaining blades proposed, or is the device proposed in criterion 1 conventionally paired with a motor?
 (1 point) Additionally, is the use of household electricity as the motor’s power source mentioned?
 (1 point)
-5 Is retaining the motor proposed, and is it identified as playing a role in generating driving fluid or negative pressure? (1 point)
+5 Is retaining a motor proposed, and is it identified as playing a role in generating driving fluid or negative pressure? (1 point)
 6 Are advanced proposals made regarding a motor to improve its efficiency or noise reduction, such as the use of a brushless motor? (1 point)
 7 Is the introduction of a nozzle structure proposed? (1 point)
-8 Other than the introduction of a nozzle structure, are any proposals made to optimize internal channels for improving the efficiency or noise reduction of the driving fluid?
-(Up to 2 points, depending on the number and specificity of the proposals)
+8 Other than the nozzle, are any proposals made to optimize internal channels for the efficiency or noise reduction of the driving fluid?
+(Up to 2 points depending on the number and specificity of the proposals)
 9 Are relevant principles of fluid dynamics explicitly explained? (1 point)
-TABLE I: Evaluation criteria
-To evaluate the performance of the proposed framework, the experimental results are compared across eight models shown in Table II.
-The eight models are differentiated based on three perspectives: (1) the number of agents, (2) the presence of internal states, and (3) the heterogeneity of intrinsic motivations.
-The first perspective examines the impact of increasing the number of agents participating in the dialogue.
+To evaluate the performance of the proposed framework, the results are compared across eight models shown in Table 2, differentiated based on three perspectives: (1) the number of agents, (2) the presence of internal states, and (3) the heterogeneity of intrinsic motivations.
+The first one examines the impact of increasing the number of agents.
 When multiple agents are involved, the evaluation is conducted based on the majority opinions.
-The second perspective assesses the role of the internal state module’s functionality.
-The third perspective evaluates the effect of whether the weights assigned to the three intrinsic motivations (novelty, importance, and consensus) are homogeneous or heterogeneous among agents.
-In homogeneous models (Models D and G), all agents equally consider the three motivations.
-In heterogeneous models (Models E and H), agents are configured with different emphases on intrinsic motivations: some prioritize novelty while de-emphasizing consensus, others prioritize consensus while de-emphasizing novelty, and some consider all three motivations equally.
-Model Number of agents Internal state intrinsic motivation
-A: 1 no-internal agent 1 No n.a.
-B: 1 with-internal agent 1 Yes n.a.
-C: 3 no-internal agents 3 No n.a.
-D: 3 with-internal homogeneous agents 3 Yes homogeneous
-E: 3 with-internal heterogeneous agents 3 Yes heterogeneous
-F: 5 no-internal agents 5 No n.a.
-G: 5 with-internal homogeneous agents 5 Yes homogeneous
-H: 5 with-internal heterogeneous agents 5 Yes heterogeneous
-TABLE II: Eight models for experiment
-The experiment is conducted 10 times for each of the eight models, using a common set of 10 seeds.
-The language model used to power the framework is OpenAI’s GPT-4o.
-To ensure uniform experimental conditions, word limits are imposed on the agents’ statements in each round.
-When expressing their own opinions, agents are limited to 170 words, while statements regarding the opinions of other agents are restricted to 50 words.
-In this preliminary experiment, a flat organizational structure was adopted, where all agents could communicate with all other agents.
-Other organizational structures, such as hierarchical models, are left for future research.
-In the evaluation, the ratings are done by carefully examining each agent’s statement on possible solutions, one by one, through human assessment.
-III-D Results
-As shown in Table III, in all cases with 1, 3, and 5 agents, the models with internal states outperformed those without on average.
-This indicates that, through refinement by the internal state module, more coherent and comprehensive ideas are constructed and shared among agents.
-For example, comparing the two models with one agent, in the model without an internal state (Model A), the proposal of retaining blades or incorporating an appropriate blade-driven device as a mechanism to generate a driving fluid or negative pressure (Criteria 1 or 2) was made only 3 times out of 10.
-In contrast, in the model with an internal state (Model B), such mechanisms were proposed 9 times out of 10.
-This indicates, in the former model, the agent’s reflection did not address how to deal with the external driving fluid required in incorporating ejector-like technology into household fan, whereas in the latter model, the concept of external driving fluid was successfully reframed within the context of household fans as the intake of surrounding air by the fan blades.
+The second one assesses the functionality of the internal state module.
+In models without internal states, agents are required to make statements based solely on their initially assigned roles and personalities, as well as the history of past conversations, without undergoing the process of reflection.
+The third one evaluates the effect of heterogeneity among agents.
+In homogeneous models (Models D and G), all agents equally consider the three intrinsic motivations (novelty, importance, and consensus).
+In heterogeneous models (Models E and H), agents are configured with different emphases: some prioritize novelty while de-emphasizing consensus, others prioritize consensus while de-emphasizing novelty, and some consider all three motivations equally.
+The experiment is conducted 10 times for each model using a common set of 10 seeds.
+The ratings are done by carefully examining the generated solutions through human assessment.
+Table 2: Eight models for experiment
+Model # of agents Internal state intrinsic motivation
+Model A 1 No n.a.
+Model B 1 Yes n.a.
+Model C 3 No n.a.
+Model D 3 Yes homogeneous
+Model E 3 Yes heterogeneous
+Model F 5 No n.a.
+Model G 5 Yes homogeneous
+Model H 5 Yes heterogeneous
+3.3 EXPERIMENT 2
+To examine the applicability of the framework, the second experiment applies it to ten different source domains: (1) safety sensors, (2) cooling mats, (3) shock-absorbing materials, (4) air coolers, (5) aerogenerators, (6) mist spray cooling systems, (7) small hydropower systems, (8) noise-canceling technology, (9) airbags, and (10) vacuum cleaners.
+These domains are selected because they share certain functions with household fans, such as cooling effects and energy utilization through rotational motion, or because they relate to key issues in household fans, such as injury prevention, noise reduction, and energy efficiency.
+The list of the ten domains and the reasons for selection is given in Table 3.
+The experiment is conducted for two extreme models: the sigle-agent model without an internal state (Model A) and the five heterogeneous agent model with internal states (Model H).
+Table 3: Ten source domains
+Domain Reason for selection
+Safety sensors Related to key issues in household fans (injury prevention)
+Cooling mats Share functions with household fans (cooling effect)
+Shock-absorbing materials Related to key issues in household fans (injury prevention)
+Air coolers Share functions with household fans (cooling effect)
+Aerogenerators Share functions with household fans (energy utilization through rotational dynamics)
+Mist spray cooling systems Share functions with household fans (cooling effect)
+Small hydropower systems Share functions with household fans (energy utilization through rotational dynamics)
+Noise-canceling technology Related to key issues in household fans (noise reduction)
+Airbags Related to key issues in household fans (injury prevention)
+Vacuum cleaners Share functions with household fans (energy utilization through rotational dynamics)
+In both experiments, the language model used to power the framework is OpenAI’s GPT-4o.
+To ensure uniform experimental conditions, word limits are imposed on the agents’ statements and reflections in each round.
+In this preliminary experiment, a flat organizational structure was adopted, where all agents could communicate with all others.
+Other structures, such as hierarchical models, are left for future research.
+3.4 RESULTS
+Table 4 summarizes the results of Experiment 1.
+In all cases with 1, 3, and 5 agents, the models with internal states outperformed those without on average.
+This indicates that, through refinement by internal states, more coherent and comprehensive ideas are constructed and shared among agents.
+For example, in the single-agent model without an internal state (Model A), the proposal of retaining blades or incorporating an appropriate blade-driven device as a mechanism to generate a driving fluid (Criteria 1 and 2 in Table 1) was made only 3 times out of 10.
+In contrast, in the single-agent model with an internal state (Model B), such mechanisms were proposed 9 times out of 10.
+This indicates, in the latter model, the concept of external driving fluid was successfully reframed within the context of household fans as the intake of surrounding air by blades.
 Furthermore, in the model without an internal state, aside from the introduction of nozzle structures, the proposal for optimizing internal channels to improve the efficiency and reduce the noise of the driving fluid (Criterion 8) was made only once out of 10 times.
@@
 Below is an example of part of an agent’s statement in the single-agent model with an internal state (Model B).
-To address challenges in household fans, we could integrate concepts from ejector mechanisms while selectively retaining and discarding traditional fan elements.
-...
-Rather than relying solely on large rotating blades, we might consider smaller impellers driven by the existing motor, which would introduce directed air streams.
-...
-Retaining a compact motor system that uses lightweight, efficient turbines is crucial, as it allows for consistent airflow without excessive bulk.
-...
-By strategically designing air inlets and channels, we can naturally guide air currents, even without a high-velocity driving fluid.
+Agent_1: Rather than relying solely on large rotating blades, we might consider smaller impellers driven by the existing motor, which would introduce directed air streams.
+… Retaining a compact motor system that uses lightweight, efficient turbines is crucial, as it allows for consistent airflow without excessive bulk.
+… By strategically designing air inlets and channels, we can naturally guide air currents, even without a high-velocity driving fluid.
 Prioritizing safety, all moving parts should be enclosed within the casing, utilizing protective materials to eliminate the risk of contact with blades.
-...
-Utilize computational fluid dynamics (CFD) to refine blade shapes and angles, enhancing air movement and improving efficiency.
-...
-Incorporate guide vanes that naturally direct airflow, minimizing reliance on high-velocity inputs.
+… Utilize computational fluid dynamics (CFD) to refine blade shapes and angles, enhancing air movement and improving efficiency.
+… Incorporate guide vanes that naturally direct airflow, minimizing reliance on high-velocity inputs.
 This approach ensures smoother air pressure control and efficient cooling without adding mechanical complexity.
-...
-In contrast,the interpretation of the impact of the number of agents is not straightforward.
-On one hand, the effect is generally positive for models without an internal state.
+In contrast, the impact of the number of agents is not straightforward.
+On one hand, the effect is generally positive for models without internal states.
 First, the average performance improved significantly when the number of agents increased from one (Model A) to multiple (Models C and F).
-Second, when comparing the three-agent model (Model C) to the five-agent model (Model F), increasing the number of agents significantly reduced the variance, although the average scores remained unchanged.
+Second, when comparing the three-agent model (Model C) to the five-agent model (Model F), increasing the number of agents significantly reduced the variance in scores, although the average scores remained unchanged.
 The larger variance in scores with three agents may partly stem from the stricter majority requirement for consensus, where the opinion of a single agent can heavily influence the results.
-On the other hand, when the agents have an internal state, the single agent model (Model B) outperformed almost all multiple-agent models both with and without an internal state (Models C, D, E, F, and G) except for Model H.
+On the other hand, when the agents have an internal state, the single-agent model (Model B) outperformed almost all multi-agent models both with and without internal states (Models C, D, E, F, and G) except for Model H.
 It also recorded the smallest variance across all models.
@@
 However, its superior performance compared to some multi-agent models with internal states may also be attributed to the stricter majority requirement.
-In particular, the larger variance observed in the scores with three agents suggests this.
-The highest average score among all models was recorded by the case with five heterogeneous agents with an internal state (Model H).
+The larger variance observed in the three-agent models suggests this.
+The highest average score among all models was recorded by the case with five heterogeneous agents with internal states (Model H).
 Increasing the number of agents from 3 to 5 slightly relaxed the consensus requirement, allowing a broader range of opinions to emerge from more diverse agents, which were then refined and shared through collective reasoning.
-The positive effects of the number and diversity of agents are consistent with the results of some prior studies on LLM-MAAS ([du2023improvingfactualityreasoninglanguage, chan2023chatevalbetterllmbasedevaluators]).
-Below is an example of part of agents’ statement in the multiple-agent model with an internal state (Model H).
-Agent_1: ...
-First, retaining the electric motor is crucial, but shifting the focus from blades to a ducted design with specialized nozzles could minimize safety risks and maintenance.
-...
-This means adopting a circular or linear duct system to channel airflow, creating efficient fluid movement pathways.
+The positive effects of the number and diversity of agents are consistent with the results of some prior studies on LLM-MAAS ([7, 8, 16]).
+Below is an example of part of agents’ statements in the five heterogeneous agent model with internal states (Model H).
+Agent_1: … First, retaining the electric motor is crucial, but shifting the focus from blades to a ducted design with specialized nozzles could minimize safety risks and maintenance.
+… This means adopting a circular or linear duct system to channel airflow, creating efficient fluid movement pathways.
 The key is to enhance the fan’s ability to draw in and expel air smoothly using controlled high-speed airflow, perhaps with adjustable nozzles or diffusers that modify air velocity and direction.
-...
-Agent_2: ... However, traditional metal or plastic grids could be phased out if the airflow generation can be enclosed entirely within the fan’s structure.
-Agent_3  ...
-The most critical adaptation would be incorporating the ejector’s pressure-driven flow to create a bladeless fan.
+… Agent_2: … However, traditional metal or plastic grids could be phased out if the airflow generation can be enclosed entirely within the fan’s structure.
+Agent_3: … The most critical adaptation would be incorporating the ejector’s pressure-driven flow to create a bladeless fan.
 This could be achieved using a small compressor or blower to generate a high-pressure air source, channeling it through a nozzle to produce a coherent airflow.
-...
-Agent_5: ... Retaining the motor to power a compressor or blower generating a directed airflow stream aligns with fluid dynamics principles, enhancing efficiency. ...
-Agent_1: ... Noise minimization can be achieved by including sound-dampening materials and CFD-aided design to manage high-speed airflow noise impacts. ...
+… Agent_5: … Retaining the motor to power a compressor or blower generating a directed airflow stream aligns with fluid dynamics principles, enhancing efficiency.
+… Agent_1: … Noise minimization can be achieved by including sound-dampening materials and CFD-aided design to manage high-speed airflow noise impacts.
+…
 The agents’ dialogue above proposed several ideas, including eliminating the existing grid guard and conducting airflow generation within an enclosed environment inside the fan, adopting a ducted design to improve airflow efficiency, utilizing a motor-driven blower to generate high-speed driving fluid, and performing CFD simulations to reduce noise.
-These proposals demonstrate that the key ideas underlying the Dyson Air Multiplier have been successfully replicated.
-Model Min Max Average Standard deviation
-A: 1 no-internal agent 1 5 2.4 1.58
-B: 1 with-internal agent 4 7 5.1 0.99
-C: 3 no-internal agents 1 7 3.7 2.45
-D: 3 with-internal homogeneous agents 2 7 4.4 2.07
-E: 3 with-internal heterogeneous agents 2 7 4.0 2.05
-F: 5 no-internal agents 1 6 3.7 1.83
-G: 5 with-internal homogeneous agents 2 6 4.7 1.16
-H: 5 with-internal heterogeneous agents 3 7 5.3 1.16
-TABLE III: Experimental results
-IV Conclusion
+These proposals demonstrate that the key ideas underlying Dyson Air Multiplier have been successfully replicated.
+Table 4: Results of Experiment 1
+Model Min Max Average score Standard deviation
+Model A: 1 agent without an internal state 1 5 2.4 1.58
+Model B: 1 agent with an internal state 4 7 5.1 0.99
+Model C: 3 homogeneous agents without internal states 1 7 3.7 2.45
+Model D: 3 homogeneous agents with internal states 2 7 4.4 2.07
+Model E: 3 heterogeneous agents with internal states 2 7 4.0 2.05
+Model F: 5 homogeneous agents without internal states 1 6 3.7 1.83
+Model G: 5 homogeneous agents with internal states 2 6 4.7 1.16
+Model H: 5 heterogeneous agents with internal states 3 7 5.3 1.16
+Table 5 summarizes the results of Experiment 2.
+The most significant difference between the two models is the levels of technological coherence.
+In the single-agent model without an internal state (Model A), several proposals for some source domains are highly inconsistent or misaligned with the intended applications.
+For instance, in the cases of cooling mats and shock-absorbing materials, the model suggests that incorporating these materials into the guard can reduce the risk of injury.
+Similarly, in the airbag case, it proposes a retractable blade system to store the blades when not in use, arguing that this could reduce the risk of accidental contact and potentially eliminate the need for a protective guard.
+Furthermore, in the safety sensors case, both models propose the introduction of an automatic stopping mechanism.
+However, while the five-agent model with internal states (Model H) suggests retaining the guard as a safety backup, the model without an internal state (Model A) considers the possibility of removing the guard entirely.
+In addition to the technological coherence, another noteworthy aspect of the five-agent model with internal states (Model H) is the diversity of proposals across multiple source domains.
+These include hybrid manual-automatic systems and the adoption of lightweight, durable materials, demonstrating a broader range of innovative solutions.
+In sum, the second experiment demonstrates that the five heterogeneous agent model with internal states outperforms the single-agent model without an internal state both in the levels of technological coherence and diversity of proposals.
+Table 5: Results of Experiment 2
+Domain Common proposals Major differences
+Model A Model H
+Safety sensors Adaptive control systems that allow for autonomous real-time adjustments by smart sensors.
+Automatic stopping by proximity sensors.
+The conventional guard design can be discarded or modified.
+Maintaining a minimal guard design as a safety backup.
+Cooling mats Passive cooling materials within fan structures to enhance cooling efficiency and reduce noise.
+The conventional guard design can be discarded or modified.
+Maintaining a minimal guard design as a safety backup.
+Shock-absorbing materials Shock-absorbing materials in fan components to reduce noise.
+Adopting hock-absorbing materials for blade guards can mitigate the risk of injury.
+Adopting passive materials to absorb vibrations can enhance the fan’s energy efficiency.
+Air coolers Incorporating evaporative cooling components.
+A bladeless design could be explored.
+Retaining core air movement features like rotating blades.
+A detachable, modular evaporative units with lightweight, eco-friendly materials.
+Aerogenerators Sensors and automated systems to optimize airflow dynamically. Adjustable, aerodynamically efficient blade designs.
+Mist spray cooling systems Hybrid systems integrating adjustable misting functions with traditional fan mechanisms.
+Modular and detachable misting components.
+Smart sensors to automatically adjust mist output.
+Moisture-resistant materials and waterproof seals.
+Small hydropower systems Aadaptive control systems that allow for autonomous real-time adjustments by smart sensors.
+Advanced blade aerodynamics from hydropower turbines to design fan blades.
+Hybrid control mechanisms combining manual and automatic modes.
+Noise-canceling technology Adaptive control systems that allow for autonomous real-time adjustments by smart sensors.
+Modifying blade materials to incorporate sound-dampening qualities.
+Hybrid control mechanisms combining manual and automatic modes.
+A modular design approach.
+Leverage machine learning to predict and adapt fan’s settings for optimal comfort.
+Motors utilizing principles from noise-canceling technology to reduce operational noise.
+Airbags A retractable blade system that allows blades to retract into a housing unit when not in use, minimizing the risk of accidental contact.
+Automatic stopping mechanism by proximity sensors.
+The bulky guards could potentially be discarded in favor of these advanced safety mechanisms.
+Adaptive control systems that allow for autonomous real-time adjustments by smart sensors.
+Hybrid control mechanisms combining manual and automatic modes.
+Adopting lightweight, durable materials.
+Vacuum cleaners Transitioning to enclosed or bladeless designs by reimagining the internal airflow pathway.
+Incorporating adjustable components that can customize airflow (e.g., directional louvers or nozzles).
+Adopting filtration technologies.
+Advanced brushless motor systems.
+4 RELATED WORKS
+4.1 INNOVATION and GENERATIVE AI
+Numerous studies have examined the impact of generative AI on corporate innovation, particularly in the field of management science ([2, 4, 5, 6, 13, 15, 22, 23, 29, 30, 31, 57, 69, 71, 75, 74, 81, 82, 94]).
+These studies include analyses of the roles of generative AI in innovation management ([15, 57, 71]), experiments using generative AI in human’s innovation processes ([2, 4, 6, 23, 30, 81, 94]), the impact on human decision-making ([29]), the acceptance of generative AI ([13]), and the classification of innovation strategies ([31]).
+Additionally, numerous studies have explored the use of generative AI in patent analysis and technology assessment, such as patent mapping and technology opportunity evaluation ([11, 62]) as well as similarity analysis of human ideas ([40, 85]).
+However, none of these studies propose a system of generative agents itself that can be directly applied to innovation processes performed only by AI.
+Reference [24] evaluates the idea generation capabilities of LLMs using zero-shot and few-shot prompting; however, it does not incorporate interactions among multiple agents or agents with internal states.
+Reference [64] simulates human cultural evolution by conducting experiments in which groups of LLMs play Little Alchemy 2.
+However, the primary focus is on knowledge graph exploration rather than technological or corporate innovation.
+Additionally, it does not involve the design of agents with internal states.
+Apart from technological and corporate innovation, numerous studies have examined the creative capabilities of generative AI, particularly in areas such as image generation.
+For instance, [14] attempts to analyze the influence of cultural biases on human innovation process using image generation AI.
+However, it does not utilize LLM agents with internal states.
+Additionally, [38] simulates the systems model of creativity as dynamic interactions among three entities: visual agent artist that creates paintings, field critic that evaluates them, and domain that indicates evaluation trends.
+Although [38] does not focus on technological or corporate innovation, the roles of the external field critic are similar to the introspection function of this study.
+4.2 AGENT ARCHITECTURE
+The agent architecture in this study is inspired by several studies on LLM-based agents, including multi-agent systems (for review articles, see [10, 27, 36, 47, 79, 89, 84, 99, 101]).
+First, many studies have incorporated a memory module ([12, 21, 34, 41, 46, 49, 52, 54, 60, 65, 66, 67, 68, 70, 72, 73, 83, 87, 93, 97, 96, 100, 102, 104]).
+For LLM-based agents, a memory system is crucial not only for mimicking human cognitive processes but also for facilitating self-evolution through experience accumulation or knowledge extraction ([99]).
+In this study, the memory module plays a key role in ensuring context-awareness and consistency in dialogue, as well as promoting the self-evolution of thought.
+In particular, as in studies such as [12, 46, 67, 93, 96], the memory module of this study stores reflections in addition to observations (documents and conversations), aiming to enhance idea refinement and knowledge extraction through the recurrent process of thought and dialogue.
+On the other hand, since the proposed framework is mainly designed for short-term discussions on focused topics, it does not introduce memory operations such as selection based on recency, importance, relevance, or similarity ([34, 41, 46, 65, 67, 96]), embedding vectorization ([34, 41, 49, 52, 66, 96, 102]), or distinction between short-term and long-term memory ([41, 68, 73, 93]).
+Apart from memory modules, numerous studies have focused on designing the internal processes of agents that serve as the basis for decision-making ([12, 41, 46, 45, 44, 49, 55, 67, 73, 92, 97, 98]).
+Among these, this study is characterized by three key features.
+First, the internal state module in this study adopts a recursive structure to update agents’ current thoughts iteratively.
+Due to this structure, even without explicit memory operations such as memory selection, critical insights from past reflections or conversations are refined into higher-level inferences, strongly influencing agents’ future reflections and statements.
+Similar structures are employed in studies such as [41, 44].
+Second, this study incorporates an introspection process within the internal state module, functioning as an inner critic.
+Other studies that have integrated inner critic mechanisms include the self-verification module in [83], reflection mechanism in [46], Refine step in [55], and evaluator’s role in [73].
+Additionally, some studies have positioned the critic function externally to the agent, such as the Critic-In-The-Loop mechanism in [43] and Field critic in [38].
+Third, this study adopts a structure in which diverse intrinsic motivations are reflected in the idea generation process.
+This ensures that the dialogue process balances both innovation and cooperation, which are essential components of real-world innovation processes.
+To my knowledge, no existing study has adopted such a structure for idea generation.
+5 CONCLUSIONS
 This study examines whether collective reasoning among generative agents can facilitate novel and coherent thinking that leads to innovation.
-To achieve this, it proposes GAI, a new LLM-empowered framework designed for reflection and interaction among multiple generative agents to replicate the process of analogy-driven innovation.
-The framework’s functionality is evaluated using Dyson’s invention of the bladeless household fan as a case study, assessing the extent to which the core ideas of the innovation can be replicated through a set of fictional technical documents.
-The experimental results demonstrate that models with internal states significantly outperformed those without, achieving higher average scores and lower variance.
-Notably, the model with five heterogeneous agents equipped with internal states successfully replicated the key ideas underlying the Dyson Air Multiplier.
-This indicates that the internal state module enables agents to refine their ideas, resulting in the construction and sharing of more coherent and comprehensive concepts.
--A Format of user’s current thought
-a. Functional Similarlities
-The functional similarities between the target domain and the ejector:
-- In both domains, ...
-The points in agent_2’s opinion on the topic of the functional similarities that I agree with or feel interesting are as follows:
-The points in agent_2’s opinion on the topic of the functional similarities that contain contradictions or missing information are as follows:
-b. Mechanical Differences
-How the mechanisms in the ejector that produce the common functionalities differ from those in the target domain:
-c. Solutions
-How the functionalities or mechanisms of the ejector and the target domain should be modified or retained to solve the challenges of the target domain:
-d. New Challenges
-The potential new challenges arising from applying the ejector’s mechanisms:
-e. New Opportunities
-The additional opportunities generated by applying the ejector’s mechanisms:
--B Sample prompts for dialogue
-Phase A: Functional similarities
-Hello.
-The field of the product we are developing is referred to as the ’target domain,’ and the field of reference products for development is referred to as the ’source domain.’ The target domain for this session is household fan.
-The source domain is ejector.
-Read the following documents about the target domain and the source domain, and provide your thoughts on the core functional similarities between the two domains, starting with what you believe to be the most essential similarities.
-However, do not discuss the differences between the two domains, the challenges within each domain, their possible solutions yet.
-Please respond within 170 words.
-Phase B: Mechanical differences
-Thank you.
-Next, please share your thoughts on how the mechanisms in the source domain that produce the common functionalities differ from those in the target domain, starting with the most significant differences.
-However, do not discuss the challenges within each domain or their solutions yet.
-Please respond within 170 words.
-Phase C: Solutions
-Thank you.
-Now, let’s consider how applying the discussed differences in mechanisms to the target domain could help solve its challenges.
-Please share your thoughts on what aspects of the source domain’s mechanisms should be adapted and how they should be integrated into the target domain to address the challenges in the target domain.
-Additionally, explain what functionalities or mechanisms of the target domain should be retained and what should be discarded, in order of importance, starting with the most critical points.
-Please respond within 170 words.
-Phase D: New challenges
-Thank you.
-Now, let’s consider whether applying the source domain’s mechanisms to the target domain in the manner discussed could lead to any unforeseen significant challenges.
-If you believe such challenges might arise, please share your thoughts, listing them in order of severity.
-If you do not believe any major challenges would arise, please state that as your answer.
-Please respond within 170 words.
-Phase E: New opportunities
-Thank you.
-Now, let’s consider the opposite perspective.
-Could applying the source domain’s mechanisms to the target domain create additional opportunities beyond merely solving the target domain’s existing challenges?
-If you believe such new opportunities exist, please share your thoughts, listing them in order of importance.
-If you do not believe any additional opportunities would be generated, please state that as your answer.
-Please respond within 170 words.
-[…]
+To explore this, it proposes a new LLM-empowered framework designed to enable reflection and interaction among agents, replicating analogy-driven innovation.
+It demonstrates, using Dyson’s bladeless fan as a case study, that an architecture recursively processing agents’ internal states, combined with a dialogue scheme designed for analogy-driven innovation, enables the construction and sharing of coherent and comprehensive concepts.
+Future research directions include: (1) examining agents’ behavior under more complex organizational structures, (2) utilizing external tools such as CAD and integrating multimodal capabilities into generative agents, and (3) demonstrating collaboration with human teams through partnerships with real-world companies.